shift summary: owl 9/22/99

John Shields (jshields@ksera)
Wed, 22 Sep 1999 07:09:49 -0500

Shift Summary for owl shift, 09/22/99

On Shift: J. Shields, A Tripathi

Hours of beam available at all intensities: 6.6
Hours of beam available at nominal intensity: 5.5
Hours of E799 data (special runs not included): 6.2
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

fixing DAQ_DEAD/live-time errors, TeV briefly down due
injection problems
========================================================================
Runs ended this shift:
Run 14664 - 12.5 M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
E799 datataking with really lousy, intermittent beam, DAQ_DEAD
errors, frontend problems, etc.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14664 ~115 01:27 "scalar live-time out of limits" and trigger
state errors. Also had "HCC busy"
Kfrending doesn't help.
Reloaded w/map-gen. Still no dice. After
more assorted fiddling, crew pages Jason L.,
who cleverly suggests power-cycling laser
repeater. Problem solved.

189 03:07 SEM too high ( 1.2E13 )

~190-198 03:08 these spills lost to a flurry of TeV aborts.

03:15 briefly no beam, due to TeV injector probs.
Beam soon returns but is intermittent, which
unfortunately continues thru the entire shift.

04:00 SEM=1.2E13 for 3 or 4 spills in a row. Crew
calls MC and asks that SEM be <= 1E13
TeV continues to have frequent aborts

04:20 TeV still having problems delivering beam at
1E13. Frequent aborts continue, beam structure
is poor. We also keep getting "DAQ_DEAD" and
live-time errors for spills w. sem > 9.5E12
(probably due to full memory). Only 50% live time.
Crew checks w. Bob T and then asks control
room to lower beam request to 8E12

04:30 SEM=8e12. 70% live, no DAQ_DEAD errors

05:48 Live-time errors rear their ugly head yet again.
As before, kfrend insufficient. Crew power-cycles
laser repeater. Again.
We are then treated to a slew of pipeline
errors. After resetting pipelines 4 or 5 times
and an equal number of kfrends, problem goes away

06:20 In the course of doing the shift checklist, crew
discovers that NM2SR is set at 2650 Amps, but
readback is only 2570 A.

06:28 tapes fill up, run is stopped. Tapes UPS006
and UP*016-018 removed

14665 1 06:39 begin new BEAM_799 run.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

NM2SR is set at 2650 Amps, readback is only 2570 A.